OpenStudy (richley):

Use the binomial theorem to express (a+ib)^n, given that {a,b} are in the domain of Real Numbers, i^2 = -1, n is in the domain of Natural Numbers, and n odd.

OpenStudy (richley):

\[(a+ib)^n\]

OpenStudy (richley):

It says to put it in the form of A+iB where A, B are real numbers expressed as summations. I've been having alot of problems with the idea of this to the nth degree.

ganeshie8 (ganeshie8):

\[(\color{blue}{a}+\color{red}{ib})^n = \sum\limits_{k=0}^n \dbinom{n}{k} \color{blue}{a}^{n-k} (\color{red}{ib})^k\] Did you get this far ?

OpenStudy (richley):

It also states I can use the following result: \[\sum_{k=0}^{n}a_{k} = \sum_{j=0}^{\frac{ n-1 }{ 2 }}(c_{2j}c_{2j+1})\]

OpenStudy (richley):

No I really am completely clueless on this topic.

OpenStudy (richley):

I struggle alot with summations.

OpenStudy (richley):

How do I get to that point?

ganeshie8 (ganeshie8):

What are \(C_{2j}, C_{2j+1}\) ? If possible could you take a screenshot of the entire question and attach please ?

OpenStudy (richley):

Its not a problem thats online. IT just says that I am allowed to use that result. The exact thing it says I can use is in that equation there.

ganeshie8 (ganeshie8):

I still don't get what are \(C_{2j}, C_{2j+1}\) ?

OpenStudy (richley):

Im not sure. I find this very difficult to comprehend in class. How did you get that result that you got earlier?

ganeshie8 (ganeshie8):

What I have earlier is the binomial theorem http://latex.artofproblemsolving.com/4/9/4/494769588adb2ac3700e25b086fe2b7d41bba70a.png

OpenStudy (richley):

And I can just plug in my results into that theorem? Let me do that and i'll post what I get.

OpenStudy (richley):

Oh I see. So I just replace all instances of a with a and all of b with ib? So ib would just be the imaginary part "b", right?

OpenStudy (richley):

How would I simplify that big equation to the form of A+iB?

ganeshie8 (ganeshie8):

Yeah, I'm trying to make sense of the result you have posted earlier. Do you have any more knowledge on that result ?

OpenStudy (richley):

I feel like the fact that i^2 = -1 is going to come in handy here, I just dont know how it helps solve this. Really sorry to use up so much of your time, I know you have other people to help too.

OpenStudy (richley):

I think that the result shown previously is the result of the summation from k=1 to n of x^n

ganeshie8 (ganeshie8):

Gotcha! You made a typo in your result

OpenStudy (richley):

Oh I did?

OpenStudy (richley):

oh is it the +?

ganeshie8 (ganeshie8):

Yes \[\sum_{k=0}^{n}a_{k} = \sum_{j=0}^{\frac{ n-1 }{ 2 }}(a_{2j} + a_{2j+1})\]

ganeshie8 (ganeshie8):

They are simply splitting the sum into two halves

OpenStudy (richley):

How does that help?

ganeshie8 (ganeshie8):

Notice the pattern i^0 = 1 i^1 = i i^2 = -1 i^3 = -i Observe that alternate powers give you real numbers 1, -1

OpenStudy (richley):

Yeah. So then i^4 would be 1 again.

OpenStudy (richley):

and i^5 would be i

OpenStudy (richley):

Do I need to find the remainder of k/4 then? Somehow?

ganeshie8 (ganeshie8):

Let's split the sum using your earlier result : \[\begin{align}(\color{blue}{a}+\color{red}{ib})^n &= \sum\limits_{k=0}^n \dbinom{n}{k} \color{blue}{a}^{n-k} (\color{red}{ib})^k\\~\\ &= \sum\limits_{j=0}^{(n-1)/2} \dbinom{n}{2j} \color{blue}{a}^{n-2j} (\color{red}{ib})^{2j} +\dbinom{n}{2j+1} \color{blue}{a}^{n-(2j+1)} (\color{red}{ib})^{2j+1} \\~\\ \end{align}\]

ganeshie8 (ganeshie8):

See if above step makes sense. I've just used your result..

OpenStudy (richley):

Yes that makes sense, so both parts of that are the two parts shown in the result, right?

ganeshie8 (ganeshie8):

Yes, if you get this far, rest should be easy

OpenStudy (richley):

Like I said, nothing about this topic is easy to me. For some reason I struggle on this particular topic alot.

ganeshie8 (ganeshie8):

Let's split the sum using your earlier result : \[\begin{align}(\color{blue}{a}+\color{red}{ib})^n &= \sum\limits_{k=0}^n \dbinom{n}{k} \color{blue}{a}^{n-k} (\color{red}{ib})^k\\~\\ &= \sum\limits_{j=0}^{(n-1)/2} \dbinom{n}{2j} \color{blue}{a}^{n-2j} (\color{red}{ib})^{2j} + \sum\limits_{j=0}^{(n-1)/2}\dbinom{n}{2j+1} \color{blue}{a}^{n-(2j+1)} (\color{red}{ib})^{2j+1} \\~\\ \end{align}\]

ganeshie8 (ganeshie8):

Look at the expressions that have `i`

OpenStudy (richley):

oh so we can get rid of the i and replace it with -1^j, right?

OpenStudy (richley):

and the second one with j+1/2

ganeshie8 (ganeshie8):

\[\large (ib)^{2j} = i^{2j}b^{2j} =(i^2)^{j}b^{2j} =?\]

OpenStudy (richley):

(-1)^jb^2j

OpenStudy (richley):

But now it depends on whether or not J is odd...

ganeshie8 (ganeshie8):

That shouldn't be a problem, all we want is to express it in the form A + iB

ganeshie8 (ganeshie8):

\[\large (ib)^{2j+1} = i^{2j+1}b^{2j+1} = i*i^{2j}b^{2j+1} =i*(i^2)^{j}b^{2j} =?\]

OpenStudy (richley):

So the i^2 = -1, right?

ganeshie8 (ganeshie8):

Yes, plugging them in we have \[\begin{align}(\color{blue}{a}+\color{red}{ib})^n &= \sum\limits_{k=0}^n \dbinom{n}{k} \color{blue}{a}^{n-k} (\color{red}{ib})^k\\~\\ &= \sum\limits_{j=0}^{(n-1)/2} \dbinom{n}{2j} \color{blue}{a}^{n-2j} (\color{red}{ib})^{2j} + \sum\limits_{j=0}^{(n-1)/2}\dbinom{n}{2j+1} \color{blue}{a}^{n-(2j+1)} (\color{red}{ib})^{2j+1} \\~\\ &= \sum\limits_{j=0}^{(n-1)/2} \dbinom{n}{2j} \color{blue}{a}^{n-2j} \color{red}{(-1)^jb^{2j}} + \sum\limits_{j=0}^{(n-1)/2}\dbinom{n}{2j+1} \color{blue}{a}^{n-(2j+1)} \color{red}{i(-1)^jb^{2j+1}} \\~\\ &= \sum\limits_{j=0}^{(n-1)/2} \dbinom{n}{2j} \color{blue}{a}^{n-2j} \color{red}{(-1)^jb^{2j}} + \color{red}{i}\sum\limits_{j=0}^{(n-1)/2}\dbinom{n}{2j+1} \color{blue}{a}^{n-(2j+1)} \color{red}{(-1)^jb^{2j+1}} \\~\\ \end{align}\]

ganeshie8 (ganeshie8):

Looks we're done ?

OpenStudy (richley):

Ohhhhh, I see, so you just move the i to the other side of the equation and it becomes A+iB where A and B are summations.... THanks a ton man!

ganeshie8 (ganeshie8):

Np :) Do you get why below splitting works ? \[\sum_{k=0}^{n}a_{k} = \sum_{j=0}^{\frac{ n-1 }{ 2 }}(a_{2j} + a_{2j+1})\]

OpenStudy (richley):

Yes

ganeshie8 (ganeshie8):

If you have time could you explain to me

OpenStudy (richley):

Thanks so much for your help man

OpenStudy (richley):

I don't unfortunately, It's getting late and I need to go to bed. I'll tag you tomorrow morning, in the meantime I gotta help this other guy out. Thanks so much for your help.

ganeshie8 (ganeshie8):

Np.. Have god sleep :)